To make ice, a freezer that is a reverse Carnot engine extracts 42 kJ as heat at -15°C during each cycle, with coefficient of performance 5.7. The room temperature is 30.3°C. How much energy per cycle is delivered as heat to the room and work per cycle is required to run the freezer?
1
Expert's answer
2020-11-17T13:03:15-0500

Solution

energy per cycle is delivered as heat to the room is

"Q_1=\frac{Q_2 T_1}{T_2}"

"Q_1=\frac{42000\times 303.5}{258}=49.4 KJ"

Work in per cycle

"W=Q_1-Q_2=49.4-42=7.4KJ"